# Leveraging Core and Uncore Frequency Scaling for Power-Efficient Serverless Workflows

### v1

@@ Abstract.
In this paper, we introduce a QoS-aware DVFS framework for serverless workflows. $\Omega$kypous regulates the end-to-end latency of serverless workflows by supplying to the system the Core/Uncore frequency combination that minimizes power consumption. With Uncore DVFS enriching the efficient power configurations space, we devise a grey-box model that accurately projects functions’ execution latency and power, to the applied Core and Uncore frequency combination. To the best of our knowledge, $\Omega$kypous is the first work that leverages Core and Uncore DVFS as an integral part of serverless workflows. Our evaluation on the analyzed Azure Trace, against state-of-the-art (SotA) power managers, demonstrates an average power consumption reduction of 9% (up to 21%) while minimizing QoS violations.
@@ 1. Introduction
Contributions. Our main contributions are: i) Detailed analysis of the combined effects—latency and power consumption—of Core and Uncore frequency scaling on serverless computing functions, alongside a comparison with the CPU-quota scaling approach. ii) We design a global gray-box model for estimating function latency that considers the dynamicity of both the Core and Uncore parts of the processor. iii) We introduce $\Omega$kypous, the first Core and Uncore frequency controller for QoS-precise serverless workflows. iv) We provide an extensive comparative evaluation study w.r.t. the respective SotA DVFS frameworks, validating the practicality and the effectiveness of $\Omega$kypous, showing its impressive capabilities in nearly eliminating QoS violations while achieving up to 21% power savings. v) In-depth characterization of $\Omega$kypous, including decomposition, noise assessment, and responsiveness analysis.
@@ 3. DVFS strategies and Serverless: Open Challenges & Opportunities
Challenge 3: Investigating functions’ performance sensitivity under Core- & Uncore-Frequency Scaling Next, we examine the impact of CFS and UFS on the latency and power consumption of serverless functions. Specifically, we deploy the thumbnailer and video processing functions from (sebs, ) and measure their tail latency and power consumption for various configurations of the Core and Uncore frequencies available in our system (as described later in §5.5). Figure 12 summarizes the results, where the green-shaded lines indicate the configured Uncore frequency level. The red line represents the system-managed UFS (default configuration), which utilizes Intel®s firmware to vary the Uncore frequency dynamically according to the CPU usage (hill2010uncore, ; gholkar2019uncore, ). Based on these outcomes, we answer two fundamental questions:
@@
1) Is direct linear proportionality between Core frequency and latency a reasonably safe assumption? Prior research (kasture2015rubik, ; gemini2020zhou, ; chou2019mudpm, ) assume linear proportionality between application latency and applied core frequency. For example, Gemini (gemini2020zhou, ) calculates latency values at different frequency levels by assuming that the product of latency and frequency remains constant. However, as demonstrated in Fig. 3a the relationship between CFS and latency varies per function. For example, in the case of video-processing, scaling the Core frequency from 1.2GHz to 2.5GHz results in $\approx$ 50% tail latency reduction while in thumbnailer it is only improved by 30%. Therefore, given the considerable variability in the frequency-to-latency relationship across different serverless functions, relying on a linear (one-to-one) proportionality is not a universally applicable assumption. However, as depicted in Fig. 1, UFS allows for leveraging iso-latency configurations for lower power consumption. Notably, in the 2.5GHz core frequency scenario, when explicitly applying UFS, power consumption is reduced by 23.5%, with the latency degrading by only 0.3%, 10% for the thumbnailer) and video-processing respectively.
@@ 5.3. Conflict Resolver
The Conflict Resolver component is designed to efficiently manage DVFS conflicts arising from differing frequency requirements among concurrently running functions. Conflicts most commonly occur within the shared Uncore components on a per-socket basis but can also emerge in scenarios where different functions are co-located on the same physical cores. To manage these conflicts, the Conflict Resolver employs two priority queues: one for Core frequency requests (CFS) and another for Uncore frequency requests (UFS). In these queues, priorities are assigned based on frequency values, with higher frequencies receiving higher priority. When the Adaptive Controller submits a new CF/UF configuration, the Conflict Resolver adds these frequency values to the corresponding priority queues. It then checks whether the highest-priority frequency in the queue is lower than any frequency currently requested by the running functions. If so, the existing frequencies are maintained to avoid disrupting the ongoing operations of these functions (Fig. LABEL:fig:conflict-resolver). However, if a new, higher-priority frequency request appears in the queue, suggesting no conflict with running functions, the Conflict Resolver then signals for these new frequencies to be applied to the system (Fig. LABEL:fig:conflict-resolver). If the queues are empty, the lowest frequencies are applied. This hierarchical approach ensures that frequency changes are appropriately adjusted, respecting the operational requirements of all active functions.
@@ 7.1. $\Omega$kypous vs power management schemes
Typical Power Management Schemes: We assess two Linux frequency governors, ondemand and conservative, excluding performance and powersave for their static frequency settings, leading to high power consumption or QoS violations. Ondemand, the default governor, exhibits inefficient power usage, while conservative, with a less aggressive frequency fluctuation, shows improved consumption. significantly enhances power efficiency, reducing consumption by an average of 16% and 11% compared to ondemand and conservative, respectively. This improvement is more pronounced in scenarios with lenient QoS requirements or the omitted (almost inactive) cluster trace.
State-of-the-art power managers: Next, we compare $\Omega$kypous with the strategies of two SotA CFS-based power managers. Gemini (gemini2020zhou, ) and ReTail (retail, ) are designed to manage distinct micro-services. To transform their implementation to the serverless workflow setup we consider two different approaches. First, we implement them considering the entire serverless workflow as a single service (ReTail, Gemini), which is managed as a whole. Second, we implement them in a decomposed manner, in which each serverless function is managed independently by a separate controller (ReTail+, Gemini+). For the latency estimations, we use the average values of the measured execution times (for each frequency level for ReTail, and the base frequency for Gemini).
@@
Power consumption: $\Omega$kypous improves power consumption up to 17% and 18% (9% and 16% on average) compared to Retail and Gemini respectively, as shown in Fig. 10b The factors that lead to this result include: i) The more fine-grained power management enabled by application decomposition into functions. ii) The introduction of UFS management which furtherly reduces power consumption as discussed in §7.2. Notably, as shown in the bottom of Fig. 10b, the $90^{th}$ percentile of power, is reduced up to 21% compared to Retail. This would allow for cluster-wide power-capping (RAPL) strategies without increasing the QoS violations.
@@ 7.2. $\Omega$kypous characterization
Impact of UF management: At next, we evaluate the additive impact of UFS management. We implement a CFS-only version of $\Omega$kypous ($\Omega$kypous-CFS), which only manages the Core frequency. In Fig. 12a, based on the Azure Trace clusters (Fig. 10a), we compare the power distributions of the two solutions across all clusters. As shown, we separate the power consumption on DRAM and PROC components. The spikes in power (in time windows equal to each workflow’s execution time) are further decreased by $\approx 15\%$. The average value is decreased by 5%/3% for DRAM, and PROC. With $\Omega$kypous-CFS, similar to $\Omega$kypous, addressing the QoS requirements, the application of strategies such as power capping in both components (DRAM, PROC) would be a better fit for the UFS-enabled version.
@@
(a) CFS-only vs. CFS/UFS
(b) Noise tolerance
@@ 8. Conclusion
In this paper, we introduced $\Omega$kypous, a novel QoS-precise DVFS framework for serverless workflows. $\Omega$kypous addresses the growing prominence of latency propagation effects in serverless architectures by employing closed-loop control to regulate end-to-end latency while minimizing power consumption through the optimized adjustment of Core and Uncore frequency combinations. A grey-box modeling approach is introduced to accurately predict execution latency and guide the coordinated CFS/UFS control decisions. We thoroughly evaluate $\Omega$kypous over the recently released Azure Trace data and we showed that $\Omega$kypous consistently outperforms state-of-the-art power managers, achieving up to 21% improved power usage on average compared to state-of-the-art power management schemes, while minimizing QoS violations and being resilient to external noise injection.

### v2

@@ Abstract.
In this paper, we introduce a QoS-aware DVFS framework for serverless workflows. $\Omega$kypous regulates the end-to-end latency of serverless workflows by supplying to the system the Core/Uncore frequency combination that minimizes power consumption. With Uncore DVFS enriching the efficient power configurations space, we devise a grey-box model that accurately projects functions’ execution latency and power, to the applied Core and Uncore frequency combination. To the best of our knowledge, $\Omega$kypous is the first work that leverages Core and Uncore DVFS as an integral part of serverless workflows. Our evaluation on the analyzed Azure Trace, against state-of-the-art (SotA) power managers, demonstrates a power consumption reduction of 20% while minimizing QoS violations.
@@ 1. Introduction
Contributions. Our main contributions are: i) Detailed analysis of the combined effects—latency and power consumption—of Core and Uncore frequency scaling on serverless computing functions, alongside a comparison with the CPU-quota scaling approach. ii) We design a global gray-box model for estimating function latency that considers the dynamicity of both the Core and Uncore parts of the processor. iii) We introduce $\Omega$kypous, the first Core and Uncore frequency controller for QoS-precise serverless workflows. iv) We provide an extensive comparative evaluation study w.r.t. the respective SotA DVFS frameworks, validating the practicality and the effectiveness of $\Omega$kypous, showing its impressive capabilities in nearly eliminating QoS violations while achieving 20% power savings. v) In-depth characterization of $\Omega$kypous, including decomposition, noise assessment, and responsiveness analysis.
@@ 3. DVFS strategies and Serverless: Open Challenges & Opportunities
Challenge 3: Investigating functions’ performance sensitivity under Core- & Uncore-Frequency Scaling Next, we examine the impact of CFS and UFS on the latency and power consumption of serverless functions. Specifically, we deploy the thumbnailer and video processing functions from (sebs, ) and measure their tail latency and power consumption for various configurations of the Core and Uncore frequencies available in our system (as described later in §5.5). Figure 12 summarizes the results, where the green-shaded lines indicate the configured Uncore frequency level. Based on these outcomes, we answer the following question:
@@
Is direct linear proportionality between Core frequency and latency a reasonably safe assumption? Prior research (kasture2015rubik, ; gemini2020zhou, ; chou2019mudpm, ) assume linear proportionality between application latency and applied core frequency. For example, Gemini (gemini2020zhou, ) calculates latency values at different frequency levels by assuming that the product of latency and frequency remains constant. However, as demonstrated in Fig. 3a the relationship between CFS and latency varies per function. For example, in the case of video-processing, scaling the Core frequency from 1.2GHz to 2.5GHz results in $\approx$ 50% tail latency reduction while in thumbnailer it is only improved by 30%. Therefore, given the considerable variability in the frequency-to-latency relationship across different serverless functions, relying on a linear (one-to-one) proportionality is not a universally applicable assumption. However, as depicted in Fig. 1, UFS allows for leveraging iso-latency configurations for lower power consumption. Notably, in the 2.5GHz core frequency scenario, when explicitly applying UFS, power consumption is reduced by 23.5%, with the latency degrading by only 0.3%, 10% for the thumbnailer) and video-processing respectively.
@@ 5.3. Conflict Resolver
The Conflict Resolver component is designed to efficiently manage DVFS conflicts arising from differing frequency requirements among concurrently running functions. Conflicts most commonly occur within the shared Uncore components on a per-socket basis but can also emerge in scenarios where different functions are co-located on the same physical cores. To manage these conflicts, the Conflict Resolver employs two priority queues: one for Core frequency requests (CFS) and another for Uncore frequency requests (UFS). In these queues, priorities are assigned based on frequency values, with higher frequencies receiving higher priority. When the Adaptive Controller submits a new CF/UF configuration, the Conflict Resolver adds these frequency values to the corresponding priority queues. It then checks whether the highest-priority frequency in the queue is lower than any frequency currently requested by the running functions. If so, the existing frequencies are maintained to avoid disrupting the ongoing operations of these functions. However, if a new, higher-priority frequency request appears in the queue, suggesting no conflict with running functions, the Conflict Resolver then signals for these new frequencies to be applied to the system. If the queues are empty, the lowest frequencies are applied. This hierarchical approach ensures that frequency changes are appropriately adjusted, respecting the operational requirements of all active functions.
@@ 7.1. $\Omega$kypous vs power management schemes
Typical Power Management Schemes: We assess two Linux frequency governors, performance and ondemand. Performance is the configuration selected to achieve the minimum execution latency. On the other hand, ondemand follows a CPU utilization-driven approach. Our solution significantly enhances power efficiency, reducing consumption by an average of 44% and 25% compared to performance and ondemand, respectively. This improvement is more pronounced in scenarios with lenient QoS requirements or the omitted (almost inactive) cluster trace.
State-of-the-art power managers: Next, we compare $\Omega$kypous with the strategies of two SotA CFS-based power managers. Gemini (gemini2020zhou, ) and ReTail (retail, ) are designed to manage distinct micro-services. We implement them considering the entire serverless workflow as a single service (ReTail, Gemini), which is managed as a whole. For the latency estimations, we use the average values of the measured execution times (for each frequency level for ReTail, and the base frequency for Gemini).
@@
Power consumption: $\Omega$kypous improves power consumption by 20% and 21% on average compared to Gemini and Retail implementations respectively, as shown in Fig. 10b The factors that lead to this result include: i) The more fine-grained power management enabled by application decomposition into functions. ii) The introduction of UFS management which furtherly reduces power consumption as discussed in §7.2. Notably, as shown in the bottom of Fig. 10b, the $90^{th}$ percentile of power, is reduced up to XXcompared to Retail. This would allow for cluster-wide power-capping (RAPL) strategies without increasing the QoS violations.
@@ 7.2. $\Omega$kypous characterization
Impact of runtime knob management: Next, we compare the impact of two different runtime configuration knobs on power efficiency: $\Omega$kypous utilizing CFS/UFS, and $\Omega$kypous-Core employing Core quotas scaling similar to (seqclock, ; zhou2022aquatope, ). Figure 12a shows the CDF of power for both approaches. As illustrated, the combined use of CFS and UFS achieves a 16% lower median power consumption than Core quotas alone. The CFS/UFS strategy also provides over 45W additional headroom for more aggressive power capping compared to Core quotas, demonstrating its significant potential.
@@
(a) Core vs. Freq. Scaling.
(b) Noise tolerance.
@@ 8. Conclusion
In this paper, we introduced $\Omega$kypous, a novel QoS-precise DVFS framework for serverless workflows. $\Omega$kypous addresses the growing prominence of latency propagation effects in serverless architectures by employing closed-loop control to regulate end-to-end latency while minimizing power consumption through the optimized adjustment of Core and Uncore frequency combinations. A grey-box modeling approach is introduced to accurately predict execution latency and guide the coordinated CFS/UFS control decisions. We thoroughly evaluate $\Omega$kypous over the recently released Azure Trace data and we showed that $\Omega$kypous consistently outperforms state-of-the-art power managers, achieving 20% improved power consumption on average, while minimizing QoS violations and being resilient to external noise injection.